OpenStudy (anonymous):

Medal Multiply or divide 5/8 (-7/12) 16/25

jagr2713 (jagr2713):

\[\frac{ 5 }{ 8 }(-\frac{ 7 }{ 12 })\times \frac{ 16 }{ 25 }\] ??

OpenStudy (anonymous):

i think so it doesnt say the sign

jagr2713 (jagr2713):

Is there choices

OpenStudy (anonymous):

no we have to determine that but i think its multiply

jagr2713 (jagr2713):

Well first we apply the rule :\[\frac{ a }{ b }\times \frac{ c }{ d }=\frac{ ac }{ bd }\]

jagr2713 (jagr2713):

So we get \[\frac{ 5\times -7\times 16 }{ 8\times 12\times 25 }\] what do we do next?

OpenStudy (anonymous):

we simplify

OpenStudy (anonymous):

Simply multiply them..ans would be with a -be sign

jagr2713 (jagr2713):

Goo so we get\[-\frac{ 560 }{ 8\times 12\times 25 }\]

jagr2713 (jagr2713):

So whats next @harz360

OpenStudy (anonymous):

-560/1500

OpenStudy (anonymous):

56/150

jagr2713 (jagr2713):

Nope we 8*12*25=2400

OpenStudy (anonymous):

-560/2400

jagr2713 (jagr2713):

Good so whats next?

OpenStudy (anonymous):

use pemdas 5/8 (-7/12) 16/25 -(5*7*16)/(25*12*8)=

OpenStudy (anonymous):

56/240

jagr2713 (jagr2713):

Next we simplify \[\frac{ 560 }{ 2400 } \ \to \frac{ 7 }{ 30 }\]

jagr2713 (jagr2713):

You get it @harz360

OpenStudy (anonymous):

what did you multiply by

OpenStudy (anonymous):

-(5*7*16)/(25*12*8)=-7/30

jagr2713 (jagr2713):

SOrry -7/30

jagr2713 (jagr2713):

@harz360 what number do you use that when you divide by 560 gives you 80 and same for 2400

jagr2713 (jagr2713):

560 is 7 right? an 2400 is 30 right?

OpenStudy (anonymous):

yes i understand

jagr2713 (jagr2713):

So we get \[-\frac{ 7 }{ 30 }\]

jagr2713 (jagr2713):

Glad to had helped :D

OpenStudy (anonymous):

thank you

jagr2713 (jagr2713):

Anytime :D